# C-3DPO: Constrained Controlled Classification for Direct Preference Optimization

### v1

@@ Abstract
Direct preference optimization (DPO)-style algorithms have emerged as a promising approach for solving the alignment problem in AI. We present a novel perspective that formulates these algorithms as implicit classification algorithms. This classification framework enables us to recover many variants of DPO-style algorithms by choosing appropriate classification labels and loss functions. We then leverage this classification framework to demonstrate that the underlying problem solved in these algorithms is under-specified, making them susceptible to probability collapse of the winner-loser responses. We address this by proposing a set of constraints designed to control the movement of probability mass between the winner and loser in the reference and target policies. Our resulting algorithm, which we call Constrained Controlled Classification DPO (C-3DPO), has a meaningful RLHF interpretation. By hedging against probability collapse, C-3DPO provides practical improvements over vanilla DPO when aligning several large language models using standard preference datasets.
Machine Learning, ICML
@@ 1 Introduction
The problem of ensuring that AI systems act in accordance with human preferences, also known as the alignment problem, has become a critical focus in machine learning. Reinforcement Learning from Human Feedback (RLHF) has emerged as a promising approach towards solving this problem (Christiano et al., 2017). RLHF proceeds by first learning a reward model (RM) from a winner-loser preference dataset followed by employing standard RL algorithms to maximize the learned RM while keeping the model close to a reference model.
@@
Recent years have witnessed the emergence of algorithms that solve the two RLHF sub-problems as a single optimization problem. Chief among these methods is the Direct Preference Optimization (DPO) algorithm (Rafailov et al., 2023). It proceeds by leveraging the closed-form solution of the RLHF objective and using the preference dataset to align the model, thus obviating the explicit reward-learning phase of RLHF. Since then, numerous extensions and successors of DPO have been proposed, highlighting the need for a deeper examination of this new class of algorithms in order to unify and connect their underlying principles, and to highlight and fix their potential pitfalls.
In this paper, we first present a classification framework that reformulates the family of DPO-style algorithms. The primary benefit of this framework is that it unifies and extends a variety of existing DPO-style algorithms. Most notably, with specific choices of classification labels and loss function, we recover the popular DPO (Rafailov et al., 2023) and IPO (Azar et al., 2024) algorithms, revealing them as special cases of our classification formulation. Moreover, unlike traditional RLHF approaches that rely solely on binary preference pairs, our classification framework can naturally incorporate richer information, such as ranked lists of preferences and auxiliary score information (e.g., ratings) about the responses.
We next leverage this classification perspective to investigate a peculiar behavior observed in several DPO-style algorithms: the decrease, or even collapse to zero, of the winner-loser probabilities. Note that prior work has documented this phenomenon (Adler et al., 2024; Xu et al., 2024; Pal et al., 2024; Fisch et al., 2024; Xiao et al., 2024; Shen et al., 2024; Wu et al., 2024; D’Oosterlinck et al., 2024). Using our proposed classification framework, we pinpoint the root cause of this behavior: minimizing the loss in the DPO-style of algorithms only provides a single constraint for learning two probabilities, leaving the problem under-constrained. Leaning on this insight, we then propose a new set of constraints that can provably fix the under-specified nature of the DPO-style algorithms, thus effectively addressing the probability-collapse issue. These constraints are designed to control the movement of winner-loser probability mass between the reference and target policies. Our proposed algorithm, Constrained Controlled Classification DPO (C-3DPO), optimizes the DPO objective under these constraints.
We evaluate the effectiveness of our set of constraint in enhancing preference alignment. Across two standard datasets and three with up to 13B parameters, C-3DPO outperforms vanilla DPO and several other baselines, delivering higher-quality final models when assessed holistically on the standard MT Benchmark (Zheng et al., 2023).
@@ 2 Preliminaries
In this section, we present the key ingredients of preference optimization, which we will build upon in the subsequent sections. In this setting, we are given a dataset $\mathcal{D}$ of triplets $(x,y_{w},y_{l})$, where $x$ is a prompt, while $y_{w}$ and $y_{l}$ reflect our preference in choosing $y_{w}$ over $y_{l}$ conditioned on $x$. We are also given a reference $\pi_{\textrm{ref}}$ serving as a guardrail.
@@
where $\beta>0$ is a hyper-parameter denoting the relative importance of reward maximization against ensuring a low deviation from $\pi_{\textrm{ref}}$. The RM is typically learned assuming that preferences follow the Bradley-Terry (BT) model:
$p(y_{w}\succ y_{l}\mid x)=\sigma\big{(}r(x,y_{w})-r(x,y_{l})\big{)}\;,$ (2)
where $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function and $r$ is the latent reward of the annotator. Therefore, fine-tuning $\pi_{\theta}$ in the RLHF approach is split into two stages: reward-learning using the BT model, followed by a policy optimization using (1).
More recently, a family of algorithms have emerged that solve the above two problems in a single stage, which is referred to as Direct Preference Optimization (DPO)-style algorithms. The key insight in DPO-style algorithms is that problem (1) admits the following closed-form solution:
$\pi^{*}(y|x)=\frac{\pi_{\textrm{ref}}(y|x)\exp\big{(}r(x,y)/\beta\big{)}}{Z(x)}\ ,$ (3)
where $Z(x)$ is the partition function and is generally intractable. We can rewrite (3) as
@@
Substituting $r(x,y)$ from (4), $Z(x)$ cancels out, leading to the optimization problem solved by DPO:
@@ 3 A Classification View of DPO-style Algorithms
In this section, we show that DPO-style algorithms can be viewed as classification algorithms. The standard classification setting has three main ingredients.
First, we typically construct a classifier hypothesis space by defining a parametric form for probabilities assigned to each class, and then train the classifier to adjust these probabilities. We show that in DPO-style algorithms these probabilities are implicitly defined as follows:
@@
where $r_{\theta}$ is the reward defined in (4) having substituted $\pi^{*}$ with $\pi_{\theta}$. It is straightforward to see that under distribution $p_{\theta}$, the probability assigned to the winner (preferred) response $y_{w}$, denoted by $p^{w}_{\theta}$, does not depend on the partition function $Z(x)$ and can be written as
@@
The probability assigned to the loser (dispreferred) response $y_{l}$, denoted by $p^{l}_{\theta}$, is defined similarly. Note that the distribution $p_{\theta}$ in (6) is a function of $\pi_{\theta}$, $\beta$, and $\pi_{\textrm{ref}}$, and can be thought of as a generalization of the conditional probability of a response $y$ given that $y\in\{y_{w},y_{l}\}$.
Second, in the standard classification setting, the dataset gives us access to labels, which we use to extract the probabilities associated with each class. To obtain these target probabilities, we simply use any distribution $p=(p^{w},p^{l})$ from the simplex $\Delta_{2}$, which is defined as the set of all vectors $p$ in $\mathbb{R}^{2}$ satisfying $p^{w},p^{l}\geq 0$ and $p^{w}+p^{l}=1$. In the most basic case we just use the one-hot vector $(p^{w},p^{l})=(+1,0)$ akin to using hard labels. This indicates that we desire to always prefer the answer $y_{w}$ to $y_{l}$ when presented with a choice between the two. More generally, we can use soft labels, meaning that we put some non-zero weight behind each class. Soft labels have shown to be quite useful in classification (Müller et al., 2019), and by extension have been useful in preference optimization where we are less certain about the given preference $y_{w}\succ y_{l}$.
Third, to solve a classification problem, we often define a differentiable loss function measuring the discrepancy between a distribution $p_{\theta}$ in the hypothesis space and the target distribution $p$. We can use any standard loss $\mathcal{L}$ for classification leading to the optimization problem: $\min_{\theta}\sum_{\mathcal{D}}{\mathcal{L}}(p_{\theta},p)$. The loss function should be non-negative $\mathcal{L}(p_{1},p_{2})\geq 0$, and satisfy $\mathcal{L}(p_{1},p_{2})=0$ if and only if $p_{1}=p_{2}$. A good example is the CE loss.
We now show that a large number of DPO-style algorithms can be viewed as specific instances of this classification framework. The generality arises from the ability to use (a) hard or soft labels for the target distribution $p$ and (b) different classification losses $\mathcal{L}$. To start we show that the popular DPO algorithm (Rafailov et al., 2023) can be recovered using hard labeling and the CE loss.
@@
Suppose that we use the CE loss and hard labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(+1,0)$ in the classification framework above. Then, using (7), we can write
${\mathcal{L}}\big{(}p_{\theta},\ p\big{)}$ $=-\left(p^{w}\log p^{w}_{\theta}+p^{l}\log p^{l}_{\theta}\right)\;\;=\;-\log p^{w}_{\theta}$
$=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\ ,$
which is exactly the DPO loss (5) summed over $\mathcal{D}$.
Another popular DPO-style algorithm is IPO (Azar et al., 2024). While the derivation of IPO in the original paper looks completely different than DPO, we now show that IPO can also be viewed as a specific instance of our classification framework. In contrast to DPO, to obtain IPO we use soft labeling in conjunction with a less common loss function which we present below.
@@
Suppose that we use the following loss
$\mathcal{L}\big{(}p_{\theta},\ p\big{)}=\big{(}\log(p^{w}_{\theta}/p^{l}_{\theta})-\log(p^{w}/p^{l})\big{)}^{2}\ ,$ (8)
and soft labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(\sigma(1/2),\sigma(-1/2))$ in the above classification framework. We will then recover the IPO loss (Eq. 17 in Azar et al. 2024). See Appendix A for more details.
Azar et al. (2024) motivated IPO primarily as a mean to overcome the issue of vanishing KL penalty in the DPO algorithm. We highlight that IPO can hedge against this issue merely because it employs soft classification labels. Stated differently, to avoid the vanishing KL, we do not have to commit to the specific IPO loss. We can achieve this desirable property using soft classification labels and for instance, the CE loss, instead of the square loss in IPO. In this case, we obtain the Conservative DPO (CDPO) algorithm (Mitchell, 2024) in which $p:=(p^{w},p^{l})=(1-\varepsilon,\varepsilon)$ for $20<\varepsilon<1/$ 2 (see more details in Appendix B). Similarly, to IPO, CDPO is free from the vanishing KL issue.
In Table 1, we show how several DPO-style algorithms can be formulated using our classification framework.
@@
IPO Soft Eq. (8)
List (Sec. 3.1) DPO(PL) Hard CE
Auxiliary Info RPO Soft CE
(Sec. 3.2) Distilled DPO Soft Eq. (8)
Table 1: Examples of DPO-style algorithms as classification. By DPO(BT) and DPO(PL), we refer to the DPO algorithm with Bradley-Terry and Plackett-Luce (see Sec. 3.1) models, respectively. More details on RPO (Adler et al., 2024) and Distilled DPO (Fisch et al., 2024) algorithms can be found in Sec. 3.2.
We now provide two straightforward extensions of the above classification framework that allow us to incorporate a larger set of existing DPO-style algorithms.
3.1 List of Preferences
Our classification framework can be extended to work with lists, rather than pairs, of preferences. In particular, assume that we have $N$ responses for each prompt $x$, giving us a dataset of the form $\mathcal{D}=\{(x,y_{1},y_{2},\ldots,y_{N})\}.$ In this case, we can define a list version of the probability vector $p_{\theta}(x,y_{1},\ldots,y_{N})$ similarly to (6), together with a target distribution $p$. With this simple extension we can now incorporate existing DPO-style algorithms that work with lists. For instance, we can show that DPO with the Plackett-Luce model for preferences (Rafailov et al., 2023) can be captured in our classification framework, again using hard labels and the CE loss. See Appendix C for a proof.
3.2 Auxiliary Information
More formally, suppose now that our dataset is comprised of ${\mathcal{D}}=(x,y_{w},s_{w},y_{l},s_{l})$. To obtain the soft labels we can employ, for instance, $p=\text{softmax}(s_{w},s_{l})$. Combining this with the IPO loss (8), we recover Distilled DPO (Eq. 7 in Fisch et al. 2024). Using the same soft labels, but with the CE loss recovers RPO (see Sec. 3.3.2 in Adler et al. 2024). We provide more details on both algorithms in Appendix D. These natural extensions further demonstrate that our classification framework is fairly general as well as sufficiently flexible to capture a large number of existing DPO-style algorithms.
4 A Pitfall of DPO-style Algorithms
We now leverage the classification framework developed in Section 3 to explain a peculiar behavior of several DPO-style algorithms, namely the decrease, or even the collapse (to 00), of the winner response probability. We then use this insight to derive algorithms that address this phenomenon in Section 5. Several prior work (Adler et al., 2024; Xu et al., 2024; Pal et al., 2024; Fisch et al., 2024; Xiao et al., 2024) reported this undesirable behavior, empirically showed that it could result in suboptimal alignment, and proposed loss functions (algorithms) to address it.
Reinspecting the classification framework allows us to clearly identify the root-cause of this behavior as the underlying problem solved in these algorithms is under-specified. In practice, the under-specified nature of the problem often manifests itself in the form of collapsing the winner-loser probabilities to 00.
To understand this phenomenon, recall from Section 3 that the original problem can be reformulated as: $\min_{\theta}\sum_{\mathcal{D}}{\mathcal{L}}(p_{\theta},p)$. For simplicity, we use the soft labels $p=(1-\varepsilon,\varepsilon)$ for some $20<\varepsilon<1/$ 2. An optimal parameter $\theta$ is one that achieves 00 loss for all samples in $\mathcal{D}$, which means that $p_{\theta}(x,y_{w},y_{l})=p$, for all $(x,y_{w},y_{l})\in\mathcal{D}$. Setting $p_{\theta}^{w}(x,y_{w},y_{l})$ in (7) equal to $p^{w}=1-\varepsilon$, we obtain
$\pi_{\theta^{*}}(y_{w}|x)=\eta\cdot\pi_{\theta^{*}}(y_{l}|x)\ ,$ (9)
with $\eta\mathrel{\mathop{:}}=\sqrt[\beta]{(1-\varepsilon)/\varepsilon}\cdot\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}$. Note that the derivation using $y_{l}$ yields the same result. Thus, we have two probabilities, $\pi_{\theta^{*}}(y_{w}|x)$ and $\pi_{\theta^{*}}(y_{l}|x)$, that we aim to learn, but minimizing the loss only gives us one constraint specified in (9). This means that the original problem is under-specified. In Figure 1, we provide an illustration, and below we give a concrete example further highlighting that the two probabilities can move in arbitrary directions.
@@ 4 A Pitfall of DPO-style Algorithms
Figure 1: An illustration of the set of solutions that achieve 00 loss in DPO-style algorithms. The shaded gray is the set of feasible solutions. The red line passing the feasible set is defined by (9) and indicates the optimal set of solutions. All points on this line achieve 00 loss. Note that the case where the probabilities collapse to 0, i.e., $p_{\theta}(x,y_{w},y_{l})=(0^{+},0^{+})$ also lies on this line.
Remark 4.1 (A Concrete Example).
Suppose we have $\varepsilon=1/11$ and for simplicity we set $\beta=1$, $\pi_{\textrm{ref}}(y_{w}|x)=0.02$, and $\pi_{\textrm{ref}}(y_{l}|x)=0.01$, which result in $\eta=20$. We identify two pairs of probabilities that satisfy (9):
  1. 1.
[…]
  2. 2.
[…]
Note that for simplicity we worked with soft labels ($\varepsilon>0$). However, notice that with hard labels we often use early stopping, which in effect prevents the model from overfitting to the hard labels. Therefore, the problem remains under-specified even when working with hard labels.
5 Constrained Controlled Classification for DPO (C-3DPO)

### v2

# C2-DPO: Constrained Controlled Direct Preference Optimization

@@ Abstract
Direct preference optimization (DPO) has emerged as a promising approach for solving the alignment problem in AI. In this paper, we make two counter-intuitive observations about DPO. First, we show that DPO loss could be derived by starting from an alternative optimization problem that only defines the KL guardrail on in-sample responses, unlike the original RLHF problem where guardrails are defined on the entire distribution. Second, we prove a surprising property of this alternative optimization problem, namely that under its optimal policy, both preferred and rejected responses tend to decrease in probability, a phenomenon typically displayed by DPO in practice. To control this behavior, we propose a set of constraints designed to limit the displacement of probability mass between the preferred and rejected responses in the reference and target policies. The resulting algorithm, which we call Constrained Controlled DPO (C2-DPO), has a meaningful RLHF interpretation. By hedging against the displacement, C2-DPO provides practical improvements over vanilla DPO when aligning several language models using standard preference datasets.
@@ 1 Introduction
Ensuring that AI systems act in accordance with human preferences, also known as the alignment problem, has become a critical focus in machine learning. Reinforcement Learning from Human Feedback (RLHF) has emerged as one promising approach (Christiano et al., 2017). RLHF proceeds by first learning a reward model (RM), and then employing standard RL algorithms to maximize the RM while keeping the model close to a reference model. Recent years have witnessed the emergence of algorithms that solve the two RLHF sub-problems in a single step, chief among them being Direct Preference Optimization (DPO) algorithm (Rafailov et al., 2023). DPO proceeds by leveraging the closed-form solution of the RLHF objective and using the preference dataset to align the model, thus obviating the explicit reward-learning, as well as the need to sample new responses during training. Since then, numerous extensions and successors have been proposed, e.g., IPO (Azar et al., 2024) and CDPO (Mitchell, 2024), underscoring the need for a deeper investigation into this emerging class of algorithms to connect the underlying principles.
In this paper, we start with a counter-intuitive observation about DPO: the DPO loss can be derived from an alternative optimization problem that imposes the KL penalty only on in-sample responses - those present in the preference dataset - rather than on the full output distribution, as done in traditional RLHF. We show that this subtle shift has a significant implication: the alternative optimization problem incentivizes in-sample probability reduction in DPO. We formally prove that under the optimal solution to this new problem, both preferred and rejected responses tend to decrease in probability. This phenomenon, while counter-intuitive, mirrors recent findings about DPO behavior (e.g., Adler et al. 2024; Xu et al. 2024; Pal et al. 2024; Fisch et al. 2024; Xiao et al. 2024; Shen et al. 2024; Wu et al. 2024; D’Oosterlinck et al. 2024), and is referred to as likelihood displacement by Razin et al. (2025). We then show that the above in-sample probability reduction phenomenon is shared among DPO extensions/successors by developing a simple classification framework that unifies the family of DPO-style algorithms.
Leaning on these insights, we propose a family of constraints that provably control likelihood displacement in DPO-style algorithms. The constraints are designed to limit the movement of winner-loser probability mass between the reference and target policies. Our proposed algorithm, Constrained Controlled DPO (C2-DPO), optimizes the DPO objective under these constraints, has a meaningful RLHF interpretation, and requires no extra computation. We evaluate the effectiveness of our constraints in enhancing preference alignment across two datasets and three models with up to 13B parameters, and show that C2-DPO outperforms vanilla DPO and several other baselines, delivering higher-quality final models when assessed holistically on the standard MT-Bench dataset (Zheng et al., 2023).
@@ 2 Preliminaries
We present the key ingredients of preference optimization on which we will build in the subsequent sections. In this setting, we are given a dataset $\mathcal{D}$ of triplets $(x,y_{w},y_{l})$, where $x$ is a prompt, while $y_{w}$ and $y_{l}$ reflect our preference in choosing response $y_{w}$ over $y_{l}$ conditioned on $x$. We are also given a reference policy $\pi_{\textrm{ref}}$ (often the SFT checkpoint $\pi_{\text{SFT}}$) which serves as a guardrail.
@@
where $\beta>0$ is a hyper-parameter denoting the relative importance of reward maximization against ensuring a low deviation from $\pi_{\textrm{ref}}$. The RM is learned by minimizing the cross-entropy (CE) loss:
$\min_{\phi}\!\!\sum_{(x,y_{w},y_{l})\in\mathcal{D}}-\log\sigma\big{(}r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\big{)}\;,$ (2)
assuming that preferences follow the Bradley-Terry (BT) model: $p(y_{w}\succ y_{l}\mid x)=\sigma\big{(}r(x,y_{w})-r(x,y_{l})\big{)}\;$ where $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function and $r$ is the latent reward of the annotator. Fine-tuning $\pi_{\theta}$ in the RLHF approach is split into two stages: reward-learning using the BT model, followed by a policy optimization using (1). More recently, a family of algorithms have emerged that solve the above two problems in a single stage: Direct Preference Optimization (DPO)-style algorithms. The loss function of DPO is derived from the RLHF problem (1) using the recipe from (Rafailov et al., 2023). The key insight here is that problem (1) admits the following closed-form solution: $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)\exp\big{(}r(x,y)/\beta\big{)}/Z(x)$ with $Z(x)$ as the partition function. We can rewrite this as
@@
Substituting $r(x,y)$ from (3) into (2), the partition function $Z(x)$ cancels out, leading to the optimization problem solved by DPO:
@@
3 KL in DPO: Implicit Guardrailing with a Counter-intuitive Side Effect
Recall from the RLHF problem (1) that large deviations from $\pi_{\textrm{ref}}$ are penalized by KL, and note that the penalty is applied to the entire distribution $\pi_{\theta}(\cdot|x)$, not only to samples from the dataset $\mathcal{D}$. One may wonder whether the KL guardrail is maintained in DPO, given the equivalence between DPO and RLHF shown in Rafailov et al. (2023).
We now show that the standard DPO loss (4) can be obtained by applying guardrails only to in-sample responses, highlighting the fact that DPO does not explicitly enforce KL beyond the data it is trained on. To this end, we replace the KL-penalty in (1) with a similar penalty, but one that only operates on in-sample responses $S_{x}=\{y_{w},y_{l}\mid(y_{w},y_{l},x)\in\mathcal{D}\}$:
$\!\!\!\!\max_{\theta}\ \mathbb{E}_{x}\Big{[}\mathbb{E}_{y\sim\pi_{\theta}}\big{[}r_{\phi}(x,y)\big{]}-\beta\!\sum_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}y\in S_{x}}}\pi_{\theta}(y|x)\log\frac{\pi_{\theta}(y|x)}{\pi_{\textrm{ref}}(y|x)}\Big{]}\ .$ (5)
When comparing to the penalty term in (5), recall that $\mathbb{KL}\big{(}\pi_{\theta}(\cdot|x)||\pi_{\textrm{ref}}(\cdot|x)\big{)}:=\sum_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}y}}\pi_{\theta}(y|x)\log\frac{\pi_{\theta}(y|x)}{\pi_{\textrm{ref}}(y|x)}$.
We now prove that starting from problem (5) and following a recipe similar to the one in Rafailov et al. (2023) described in Section 2, we ultimately arrive at the same standard DPO loss (4). We report the proof of the following lemma in Appendix A.
Lemma 3.1.
Problem (5) has a closed-form solution. While this closed-form solution is different than the closed-form solution to problem (1), substituting it into the BT model and following the recipe of Rafailov et al. (2023) leads to the same standard DPO loss in (4).
At first glance, the two optimization problems (1) and (5) look quite similar, but finding the closed-form solution of (5) requires a delicate analysis. In particular, the new penalty is only summing over the set $S_{x}$, so the resultant KKT optimality conditions are more involved (details in Appendix A).
Note that the introduced penalty in (5) is not even a proper divergence, in the sense that it could be negative. Thus, the lemma could be viewed as evidence that when we move to DPO, we no longer enforce the $\mathbb{KL}$ penalty explicitly. However, empirically we still see that DPO can effectively guardrail using KL since in practice, using larger values of $\beta$ results in smaller KL deviations during DPO training (see Figure 1-Left). This arguably makes sense, because in DPO there is little incentive for the model to shift probability mass for responses that are quite different from in-sample responses, and so explicit out-of-sample guardrailing may not be necessary.
Figure 1: Left: Increasing $\beta$ in DPO leads to effective out-of-sample guardrailing. We ran DPO with different values of $\beta$ starting from the Zephyr-7B initial checkpoint on the UltraFeedback dataset. We then estimated the KL divergence between $\pi_{\theta}$ and $\pi_{\textrm{ref}}$ by autoregressively sampling $N=32$ responses from $\pi_{\theta}$ for each prompt in the test set, followed by computing $\frac{1}{N}\sum_{i=1}^{N}\log\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\textrm{ref}}(y_{i}|x)}$ and averaging over prompts. Right: Reduction of in-sample probabilities in DPO training.
While this minimal guardrailing is effective, we show that it nevertheless leads to counter-intuitive behavior. Notice again that the new penalty term in (5) can be negative, in sharp contrast to the original KL term, which is non-negative by definition. More importantly, the overall objective in (5) can be increased by decreasing the probability of both winner and loser responses, making the new penalty term negative, i.e., $\log\pi_{\theta}(y\mid x)/\pi_{\text{ref}}(y\mid x)<0,\;\forall y\in S_{x}$. Note that this behavior is not incentivized in the original optimization problem (1). We know that the KL penalty in (1) is always non-negative, so even if we reduce the in-sample portion of the KL by decreasing in-sample probabilities, the out-of sample portion must get more and more positive, and so there is no point in blindly reducing the in-sample probabilities. We now formalize our intuitions:
Lemma 3.2.
Let $(x,y)$ be an in-sample prompt-response pair, i.e., $x\in\mathcal{D}$ and $y\in S_{x}$. Suppose that $r_{\phi}(x,y)\leq\max_{y^{\prime}\notin S_{x}}r_{\phi}(x,y^{\prime})$. Then, any optimal solution $\theta$ to the optimization problem (5) satisfies $\pi_{\theta}(y|x)\leq e^{-1}\pi_{\textrm{ref}}(y|x)$.
The proof, reported in Appendix A, hinges on the fact that the penalty term in (5) is defined on in-sample responses. Therefore, this result does not hold when solving problem (1), indicating that the standard RLHF formulation is not susceptible to this counter-intuitive behavior. To better understand the result, note that for a prompt $x$, if the reward of an in-sample response (whether the response is preferred or rejected) is smaller than the maximal reward of out-of-sample responses, then any optimal solution of (5) decreases the probability of this in-sample response. Clearly the size of the in-sample response set $S_{x}$ is relatively small in comparison to the rest of the set (in the extreme case, a single preferred and rejected response), so the condition is likely to hold.
Interestingly, it has been observed recently that during DPO training all in-sample probabilities - even those associated with preferred responses - tend to decrease in magnitude (e.g., Adler et al. 2024; Xu et al. 2024; Pal et al. 2024; Fisch et al. 2024; Xiao et al. 2024; Liu et al. 2024; Yin et al. 2024; Guo et al. 2024; Yuzi et al. 2025; Razin et al. 2025; Deng et al. 2025; Xiliang et al. 2025; Huang et al. 2025; Yunan et al. 2025). We also observe this clear trend in our experiments, as is apparent in Figure 1 (Right) where we show that the sum of log ratios ($\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}+\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}$) tends to decrease radically. Lemma 3.2 hints at the underlying reason for this counter-intuitive behavior, which has been referred to as likelihood displacement of in-sample probabilities (Razin et al., 2025). To the best of our knowledge, while this has been reported in previous empirical studies, Lemma 3.2 is among very few theoretical results explaining this counter-intuitive phenomenon.
We conclude this section by noting that we proved in-sample probability reduction for DPO alone, and so it would be natural to ask if some of its main successors, e.g., IPO (Azar et al., 2024) and CDPO (Mitchell, 2024), share the same property. In the next section, we answer this question affirmatively by developing a classification framework that unifies the family of DPO-style algorithms.
@@ 4 A Classification View of DPO-style Algorithms
We now show that DPO-style algorithms can be interpreted as classification methods, where the objective is defined solely over in-sample responses. As a result, similar to DPO, none of these algorithms applies any direct guardrailing on out-of-sample responses. Recall that the standard classification setting has three main ingredients.
First, we construct a hypothesis space by defining probabilities assigned to each class. In DPO-style algorithms, these probabilities are implicitly defined as:
@@
where $r_{\theta}$ is the reward defined in (3) having substituted $\pi^{*}$ with $\pi_{\theta}$. Under $p_{\theta}$, the probability assigned to the winner (preferred) response $y_{w}$, denoted by $p^{w}_{\theta}$, does not depend on the partition function $Z(x)$ and can be written as:
@@
The probability assigned to the loser (rejected) response $y_{l}$, denoted by $p^{l}_{\theta}$, can be defined similarly. Note that the distribution $p_{\theta}$ in (6) can be thought of as a generalization of the conditional probability of a response $y$ given that $y\in\{y_{w},y_{l}\}$.
Second, in the standard classification setting, the dataset gives us access to labels, which we use to extract target probabilities associated with each class. To obtain these target probabilities, we simply use any distribution $p=(p^{w},p^{l})$ from the simplex $\Delta_{2}$, which is defined as the set of all vectors $p\in\mathbb{R}^{2}$ satisfying $p^{w},p^{l}\geq 0$ and $p^{w}+p^{l}=1$. In the most basic case, we just use the one-hot vector $(p^{w},p^{l})=(+1,0)$ akin to using hard labels. More generally, we can use soft labels, meaning that we put some non-zero weight behind each class (Müller et al., 2019).
Third, we define a classification loss $\mathcal{L}$ between two distributions $p_{\theta}$ and $p$, leading us to the optimization problem: $\min_{\theta}\sum_{\mathcal{D}}{\mathcal{L}}(p_{\theta},p)$. A good example is the CE loss.
We can now show that a large number of DPO-style algorithms can be viewed as specific instances of this classification framework. The generality arises from the ability to use (a) hard or soft labels for the target distribution $p$ and (b) different classification losses $\mathcal{L}$.
@@
Suppose that we use the CE loss and hard labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(+1,0)$ in the above classification framework. Then, using (7), we can write
${\mathcal{L}}\big{(}p_{\theta},\ p\big{)}\!=\!-\left(p^{w}\log p^{w}_{\theta}+p^{l}\log p^{l}_{\theta}\right)\!=\!-\log p^{w}_{\theta}={\displaystyle-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)}\ ,$
which is exactly the DPO loss (4) if it is summed over $\mathcal{D}$.
Another popular DPO-style algorithm is IPO (Azar et al., 2024). While the derivation of IPO in the original paper looks completely different than DPO, we now show that IPO can also be viewed as a specific instance of our classification framework (see Appendix B.1 for the detailed derivation).
@@
We can recover IPO (Eq. 17 in Azar et al. 2024) using the loss $\mathcal{L}\big{(}p_{\theta},\ p\big{)}=\big{(}\log(p^{w}_{\theta}/p^{l}_{\theta})-\log(p^{w}/p^{l})\big{)}^{2}$ and soft labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(\sigma(1/2),\sigma(-1/2))$.
Table 1 shows that several DPO-style algorithms can be formulated using this framework. Given this framework, we can formulate the set of optimal solutions (those that achieve 0 loss) for any DPO-style algorithm. An optimal parameter $\theta$ is one that achieves 00 loss for all samples in $\mathcal{D}$, i.e., $p_{\theta}(x,y_{w},y_{l})=p,\forall(x,y_{w},y_{l})\in\mathcal{D}$. Setting $p_{\theta}^{w}(x,y_{w},y_{l})$ in (7) equal to $p^{w}=1-\varepsilon$, we obtain
$\pi_{\theta^{*}}(y_{w}|x)=\eta\cdot\pi_{\theta^{*}}(y_{l}|x),\quad\;\textrm{with}\;\;\;\eta\mathrel{\mathop{:}}=\sqrt[\beta]{(1-\varepsilon)/\varepsilon}\cdot\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\ .$ (8)
Note that the derivation using $y_{l}$ yields the same result. Thus, we have two probabilities, $\pi_{\theta^{*}}(y_{w}|x)$ and $\pi_{\theta^{*}}(y_{l}|x)$, that we aim to learn, but minimizing the loss only gives us one constraint specified in (8). This means that the original classification problem (loss-minimization in DPO-style algorithms) is under-specified. In Figure 2, we provide an illustration of this phenomenon. We also give a concrete example in Appendix C in which the two probabilities can move in arbitrary directions, and most notably, they can both go to zero.
@@
IPO Soft Remark 4.2
List (Appendix. B) DPO(PL) Hard CE
Auxiliary Info RPO Soft Appendix B
(Appendix. B) Distilled DPO Soft Appendix B
Table 1: Unifying DPO-style algorithms. By DPO(PL), we mean DPO with Plackett-Luce model. See our proofs for RPO (Adler et al., 2024) and Distilled DPO (Fisch et al., 2024) in Appendix B.
@@
Figure 2: An illustration of the set of solutions that achieve 00 loss in DPO-style algorithms. The shaded gray is the set of feasible solutions. The red line passing through the feasible set indicates the set of optimal solutions. Note that the case where the probability belonging to in-sample responses displace entirely to out-of-sample responses, i.e., $p_{\theta}(x,y_{w},y_{l})=(0^{+},0^{+})$ also lies on this line.
5 Constrained Controlled DPO (C2-DPO)